OpenStudy (anonymous):

Which expression is not equivalent to the other expressions? -6(2x - 4) -(12x - 6) + 18 -3(4x - 3) + 15 -4(3x + 6)

ganeshie8 (ganeshie8):

looks in every expression has : \(-12x\)

ganeshie8 (ganeshie8):

so the difference will be in constant term

ganeshie8 (ganeshie8):

simplify each of the expression and find out the odd one

ganeshie8 (ganeshie8):

-6(2x - 4) = -12x + 24 -(12x - 6) + 18 = -12x +6 + 18 = -12x + 24 -3(4x - 3) + 15 = ? -4(3x + 6) = ?